# Energy-Conserving Data-Driven Discretizations for Maxwell’s Equations

### v1

@@ 1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [taf-love]. Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions. Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, e.g., via regression or neural networks. Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [bar-sinai].
@@ 1.1 Related work
There is a growing literature on learning discretizations and coarse-grid operators for PDEs directly from data. A prominent example is the work of [bar-sinai], who learn finite-difference schemes for Burgers and Navier–Stokes equations by fitting local stencils to high-fidelity solutions while optimizing accuracy on a training set. Related approaches use regression or neural networks to construct reduced models or learned fluxes for time-dependent PDEs; see, e.g., [parish2016paradigm, ling2016reynolds, rudy2017data, raissi2018hidden, kochkov2021machine]. In most of these methods, the learned operators are chosen solely to minimize an empirical error measure and are not required to satisfy structural constraints such as conservation, stability, or symmetries of the underlying PDE. In contrast, our focus is on learning a discrete Maxwell operator that provably satisfies a semi-discrete energy identity through an explicit skew-adjointness constraint on the convolution stencil.
@@
Classical finite-difference time-domain (FDTD) methods for Maxwell’s equations, such as the Yee scheme and its variants, are carefully constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete or fully discrete level [taf-love, richtmyer1967, morton1994, leveque2007]. In particular, staggered spatial discretizations and symplectic or A-stable time integrators such as the Crank–Nicolson scheme are often used to obtain skew-adjoint or nearly skew-adjoint update operators, leading to strong stability and conservation properties [crank1947, giles2010, hairerwanner2010]. There is also a large body of work on structure-preserving and geometric integrators for Hamiltonian ODE/PDE systems, which emphasizes invariants such as energy and symplectic form [hairer]. Our work is closest in spirit to these structure-preserving schemes, but differs in that the discrete spatial operator is learned from data rather than hand-designed: we encode skew-adjointness as a set of linear constraints on the stencil coefficients and solve a convex quadratic program to obtain an energy-conserving discretization.
@@ 2.1 Skew-symmetric convolution operators and stencil constraints
where $j=i+\ell$ (modulo $N$). Hence $D_{ij}=w_{i-j}$. ∎
@@
Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$. Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
@@ 3 Wave speed and CFL for the learned operator
If one were to apply the classical leapfrog scheme to the semi-discrete system, stability of the scalar test equation $\dot{u}=\lambda u$ with $\lambda=\mathrm{i}\omega$ requires $\Delta t\,|\omega|\leq 2$, see, e.g., [richtmyer1967, morton1994, leveque2007]. Substituting $\omega=\pm|\mu(\theta)|$ and maximizing over $\theta$ yields the CFL condition
@@ 3.1 Crank–Nicolson time integration
is preserved exactly by (15) for any time step $\Delta t>0$; see, e.g., [hairer, hairerwanner2010] for background on such structure-preserving schemes.
@@ 4.1 Training data from spectral 1D Maxwell
We denote by $U\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$ the array of training states and by $dU/dt$ the corresponding array of time derivatives, with channel 0 equal to $E$ and channel $1$ equal to $H$.
@@ 4.2 Least-squares formulation
where $\lambda\geq 0$ is a small Tikhonov regularization parameter. The function build_Ab_from_data in the code constructs $A$ and $b$.
@@ 6.1 Experimental setup
  • Time integration: Crank–Nicolson (CN) applied to the semi-discrete system with time step $\Delta t=0.5\Delta x$ and $N_{t}=300$ time steps. For skew-adjoint $D$, the CN update is norm-preserving, so no CFL restriction is required for stability, and the discrete energy is conserved up to roundoff.
@@
$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1).$
@@ 6.2 Learned stencils for $R=1$
For $R=1$ the true centered-difference stencil (in normalized form) is $w_{\mathrm{exact}}=(-1,0,1)/(2\Delta x)$. Table 1 shows the learned stencils obtained by each solver compared with $w_{\mathrm{exact}}$. We observe that all methods recover a stencil very close to the exact one, with small differences due to the finite training set and regularization.
@@
Table 1: Learned stencils for radius $R=1$. The exact centered difference (CD) stencil is compared against those learned by PG, NAG, ADMM, and CVX (values shown here are representative; see the accompanying code for the precise output on a given run).
Table 1 shows that all solvers recover a stencil very close to the exact centered difference. In Figure 1 we compare the final-time electric field $E(x,T)$ for each learned stencil against the reference FD solution; the curves lie almost on top of each other.
@@
Figure 2: Discrete energy error $E^{n}-E^{0}$ for the Crank–Nicolson time-stepping applied to the semi-discrete Maxwell system with a skew-symmetric discrete derivative $D$ (FD stencil, $R=1$). The error remains at the level of $10^{-14}$ over $N_{t}=300$ steps, consistent with conservation up to machine precision.
Figure 2 shows the discrete energy error $E^{n}-E^{0}$ over $N_{t}=300$ Crank–Nicolson steps. The drift remains at the level of $10^{-14}$, confirming that the fully discrete scheme preserves the semi-discrete energy up to roundoff.
@@ 6.3 Error and energy residual versus stencil radius
Figure 4 reports the energy-constraint residual $r_{\mathrm{eq}}(w)=\lVert Cw-d\rVert_{2}$ as a function of the stencil radius $R$. The projected-gradient and Nesterov methods enforce the linear constraints to essentially machine precision for all radii, with residuals near $10^{-16}$, since each iteration includes an explicit orthogonal projection onto the affine subspace $\{w:Cw=d\}$. ADMM also attains very small residuals, in the range $10^{-15}$$10^{-13}$. The CVX solutions exhibit somewhat larger residuals, around $10^{-9}$$10^{-7}$, but these are still tiny on the scale of the stencil coefficients (which are $\mathcal{O}(10^{1})$). Thus all methods produce stencils that are effectively skew-symmetric, but ADMM and CVX combine this with significantly smaller PDE errors in Figure 3. In particular, these results show that driving $r_{\mathrm{eq}}(w)$ all the way to machine precision (as PG and NAG do) is not sufficient for accuracy if the least-squares objective is not simultaneously minimized; ADMM and CVX achieve a more favorable balance between data fit and constraint enforcement.
Figure 4: Energy-constraint residual $\|Cw-d\|_{2}$ versus stencil radius $R$ for each solver.
@@ 6.4 Solver convergence and runtime
Finally, we compare the convergence histories of the different solvers. Figure 5 plots the objective value versus iteration for each first-order method for a fixed radius. Figure 6 shows the same objective versus wall-clock time. These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
Figure 5: Objective value versus iteration for PG, NAG, and ADMM for a fixed radius $R$.
Figure 6: Objective value versus runtime for PG, NAG, and ADMM for a fixed radius $R$.
Figure 5 shows that all three methods reduce the objective by several orders of magnitude, but their convergence patterns differ. PG exhibits the expected monotone, roughly geometric decrease. NAG achieves a much faster reduction in the first $\sim 20$ iterations, but with the characteristic oscillations of accelerated gradient methods before settling near the optimum. ADMM appears as an almost flat curve at the bottom of the plot because it solves a small KKT system at each iteration; in this setting, one or two ADMM iterations are enough to reach an objective value comparable to the final PG/NAG iterates.
Figure 6 plots the same objective values against wall-clock time. Because PG and NAG have very cheap iterations, their curves almost coincide and reach a low objective well below $10^{-3}$ in about $10^{-2}$ seconds for our problem size. ADMM has a higher per-iteration cost due to the linear solve, but it reaches essentially its final objective in a single step, so the total runtime is competitive with the first-order methods in this regime. For larger-scale problems, we expect the trade-off between cheap but many iterations (PG/NAG) and expensive but very few iterations (ADMM) to be even more pronounced.
@@ 7 Conclusion
We proposed an energy-stable, data-driven discretization framework for the one-dimensional Maxwell system. By learning convolution stencils from spectral training data under a linear skew-symmetry constraint, we obtain discrete derivative operators that are guaranteed to be skew-adjoint and therefore exactly conserve semi-discrete electromagnetic energy. We demonstrated through numerical experiments that the resulting schemes match the accuracy of standard central differences while preserving structural properties. Using Crank–Nicolson time integration, we further confirmed that the discrete energy remains constant to roundoff for both hand-crafted and learned stencils satisfying the skew-adjoint constraint.
Although the present work focuses on a simple 1D model, the ideas are readily extensible to higher dimensions and more complex material models. Our results suggest that combining data-driven discretization with structure-preserving constraints is a promising direction for building stable, physically faithful reduced models in computational electromagnetics and beyond.
@@ 8 Discussion and extensions
Future work will learn staggered curl operators that preserve a discrete Poynting theorem in 2D/3D, integrate material models (Lorentz/Duffing) with structure-preserving polarization updates, and incorporate boundary treatments (ABC/CPML). On the optimization side, extending the comparison to higher-order stencils, non-quadratic objectives, and learned nonlinear fluxes would help clarify the regimes in which ADMM or interior-point methods offer the best tradeoff relative to accelerated first-order algorithms. Ultimately, combining learning with physics-motivated constraints promises accurate, stable reduced models for large-scale CEM while keeping the mathematical guarantees that make FDTD dependable [taf-love, hairer].

### v2

@@ 1 Introduction
Finite-difference time-domain (FDTD) methods are a standard tool for solving Maxwell’s equations in computational electromagnetics [15]. Classical schemes, such as the Yee method, are derived by hand from the continuous equations, with spatial derivatives approximated using fixed central-difference stencils and stability enforced via analytical CFL conditions. Recently, there has been growing interest in data-driven discretizations, where discrete operators are learned from numerical or experimental data, e.g., via regression or neural networks. Examples include data-driven discretization of Burgers and Navier–Stokes equations and learned coarse-grid operators for PDEs [1].
@@ 1.1 Related work
There is a growing literature on learning discretizations and coarse-grid operators for PDEs directly from data. A prominent example is the work of [1], who learn finite-difference schemes for Burgers and Navier–Stokes equations by fitting local stencils to high-fidelity solutions while optimizing accuracy on a training set. Related approaches use regression or neural networks to construct reduced models or learned fluxes for time-dependent PDEs; see, e.g., [11, 9, 14, 12, 7]. In most of these methods, the learned operators are chosen solely to minimize an empirical error measure and are not required to satisfy structural constraints such as conservation, stability, or symmetries of the underlying PDE. In contrast, our focus is on learning a discrete Maxwell operator that provably satisfies a semi-discrete energy identity through an explicit skew-adjointness constraint on the convolution stencil.
@@
Classical finite-difference time-domain (FDTD) methods for Maxwell’s equations, such as the Yee scheme and its variants, are carefully constructed so that discrete versions of divergence constraints and energy identities hold at the semi-discrete or fully discrete level [15, 13, 10, 8]. In particular, staggered spatial discretizations and symplectic or A-stable time integrators such as the Crank–Nicolson scheme are often used to obtain skew-adjoint or nearly skew-adjoint update operators, leading to strong stability and conservation properties [3, 4, 6]. There is also a large body of work on structure-preserving and geometric integrators for Hamiltonian ODE/PDE systems, which emphasizes invariants such as energy and symplectic form [5]. Our work is closest in spirit to these structure-preserving schemes, but differs in that the discrete spatial operator is learned from data rather than hand-designed: we encode skew-adjointness as a set of linear constraints on the stencil coefficients and solve a convex quadratic program to obtain an energy-conserving discretization.
@@ 2.1 Skew-symmetric convolution operators and stencil constraints
where $j=i+\ell$ (modulo NN). Hence $D_{ij}=w_{i-j}$.[2]
@@
[2]Let $D$ be the convolution operator (2) with periodic boundary conditions and stencil $w\in\mathbb{R}^{2R+1}$. Then $D$ is skew-adjoint with respect to (4), i.e., $D^{\top}=-D$, if and only if the stencil satisfies
@@ 3 Wave speed and CFL for the learned operator
If one were to apply the classical leapfrog scheme to the semi-discrete system, stability of the scalar test equation $\dot{u}=\lambda u$ with $\lambda=\mathrm{i}\omega$ requires $\Delta t\,|\omega|\leq 2$, see, e.g., [13, 10, 8]. Substituting $\omega=\pm|\mu(\theta)|$ and maximizing over $\theta$ yields the CFL condition
@@ 3.1 Crank–Nicolson time integration
is preserved exactly by (15) for any time step $\Delta t>0$; see, e.g., [5, 6] for background on such structure-preserving schemes.
@@ 4.1 Training data from spectral 1D Maxwell
We denote by
$U\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$
the array of training states, where $U[s,0,i]=E^{(s)}(x_{i})$ and $U[s,1,i]=H^{(s)}(x_{i})$ denote the electric and magnetic fields of the $s$-th training sample evaluated at the grid point $x_{i}$. Similarly,
$\frac{dU}{dt}\in\mathbb{R}^{n_{\mathrm{sims}}\times 2\times N}$
stores the corresponding time derivatives, with $(dU/dt)[s,0,i]=\partial_{t}E^{(s)}(x_{i})$ and $(dU/dt)[s,1,i]=\partial_{t}H^{(s)}(x_{i})$.
Here $n_{\mathrm{sims}}$ denotes the number of independent training samples (or snapshots), each consisting of a randomly generated Maxwell state $(E,H)$ and its corresponding time derivative. The spatial domain $[0,L]$ is discretized using $N$ uniformly spaced grid points $x_{i}=i\Delta x$, $i=0,\dots,N-1$, where $\Delta x=L/N$.
@@ 4.2 Least-squares formulation
where $\lambda\geq 0$ is a small Tikhonov regularization parameter.
@@ 6.1 Experimental setup
  • Time integration: Crank–Nicolson (CN) applied to the semi-discrete system with time step $\Delta t=0.5\Delta x$ and $N_{t}=300$ time steps.
@@
$w_{\mathrm{exact}}=\frac{1}{2\Delta x}(-1,\,0,\,1),$ (22)
@@ 6.2 Learned stencils for $R=1$
For $R=1$ the true centered-difference stencil is precisely $w_{\mathrm{exact}}$ from (22).
Table 1 shows the learned stencils obtained by each solver compared with $w_{\mathrm{exact}}$. We observe that all methods recover a stencil very close to the exact one, with small differences due to the finite training set and regularization.
@@
Table 1: Learned stencils for radius $R=1$.
Table 1 shows that all solvers recover a stencil very close to the exact centered difference. In In Figures 1 and 2 we compare the space–time evolution and the final-time profile of the electric field $E(x,t)$ for each learned stencil against the reference FD solution. The space–time plots in Figure 1 show that the learned stencils reproduce essentially the same wave propagation as the exact centered difference, while the final-time profiles in Figure 2 are visually indistinguishable from the reference solution.
Figure 1: Space–time evolution of the electric field $E(x,t)$ for the reference central-difference stencil and the learned stencils with radius $R=1$. Each panel shows $E(x,t)$ over the full simulation time.
@@ 6.3 Error and energy residual versus stencil radius
6.4 Discrete energy conservation for learned Maxwell stencils
We next examine how closely the learned stencils preserve the semi-discrete electromagnetic energy under Crank–Nicolson time integration.
Figure 4: Discrete energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ for Crank–Nicolson time stepping with the exact central-difference stencil and the structure-preserving learned stencils, shown on a linear (left) and logarithmic (right) scale.
Figure 4 compares the discrete electromagnetic energy error $\mathcal{E}^{n}-\mathcal{E}^{0}$ for Crank–Nicolson time integration using the exact central-difference stencil and the structure-preserving learned stencils. On the linear scale, all curves are essentially flat and remain at the level of $10^{-12}$, indicating that the total energy is conserved up to machine precision for all methods. The logarithmic view reveals small differences: the PG, ADMM, and CVX stencils produce energy drift of the same order as the exact FD stencil, while NAG exhibits a slightly larger but still very small drift. These results confirm that enforcing the skew-symmetry constraints during training yields learned stencils that are effectively energy-preserving under Crank–Nicolson time stepping.
@@ 6.5 Solver convergence and runtime
Finally, we compare the convergence histories of the different solvers. Figure 5(a) plots the objective value versus iteration for each first-order method for a fixed radius. Figure 5(b) shows the same objective versus wall-clock time. These plots illustrate the tradeoffs among PG, NAG, and ADMM in terms of iterations, per-iteration cost, and constraint enforcement.
(a) Objective vs iteration
(b) Objective vs runtime
Figure 5: Convergence behavior of PG, NAG, and ADMM for a fixed stencil radius $R$. Left: objective value versus iteration count. Right: objective value versus wall-clock runtime.
Figure 5(a) shows that all three methods reduce the objective by several orders of magnitude, but their convergence patterns differ. PG exhibits the expected monotone, roughly geometric decrease. NAG achieves a much faster reduction in the first $\sim 20$ iterations, but with the characteristic oscillations of accelerated gradient methods before settling near the optimum. ADMM appears as an almost flat curve at the bottom of the plot because it solves a small KKT system at each iteration; in this setting, one or two ADMM iterations are enough to reach an objective value comparable to the final PG/NAG iterates.
Figure 5(b) plots the same objective values against wall-clock time. Because PG and NAG have very cheap iterations, their curves almost coincide and reach a low objective well below $10^{-3}$ in about $10^{-2}$ seconds for our problem size. ADMM has a higher per-iteration cost due to the linear solve, but it reaches essentially its final objective in a single step, so the total runtime is competitive with the first-order methods in this regime. For larger-scale problems, we expect the trade-off between cheap but many iterations (PG/NAG) and expensive but very few iterations (ADMM) to be even more pronounced.
@@ 7 Conclusion
We proposed an energy-conserving, data-driven discretization framework for the one-dimensional Maxwell system. By learning convolution stencils from spectral training data under a linear skew-symmetry constraint, we obtain discrete derivative operators that are guaranteed to be skew-adjoint and therefore exactly conserve semi-discrete electromagnetic energy. Numerical experiments show that the resulting schemes match the accuracy of standard central differences while preserving these structural properties. Using Crank–Nicolson time integration, we further confirmed that the discrete energy remains constant to roundoff for learned stencils satisfying the skew-adjoint constraint.
@@ 8 Discussion and extensions
Future work will learn staggered curl operators that preserve a discrete Poynting theorem in 2D/3D, integrate material models (Lorentz/Duffing) with structure-preserving polarization updates, and incorporate boundary treatments (ABC/CPML). On the optimization side, extending the comparison to higher-order stencils, non-quadratic objectives, and learned nonlinear fluxes would help clarify the regimes in which ADMM or interior-point methods offer the best tradeoff relative to accelerated first-order algorithms. Ultimately, combining learning with physics-motivated constraints promises accurate, stable reduced models for large-scale CEM while keeping the mathematical guarantees that make FDTD dependable [15, 5].